# Tradeoffs of Diagonal Fisher Information Matrix Estimators

### v2

@@ 5.2 Classification: Categorical Distribution
Suppose that the exponential family in Eq. 1 is a categorical distribution. With $\sigma_{\max}(\bm{x})=\max_{k}\sigma_{k}(\bm{x})$:
$\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))\leq m(\bm{x}),$
$\lambda_{\max}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}(\bm{h}\,|\,\bm{x})\otimes\mathcal{I}(\bm{h}\,|\,\bm{x}))\leq 2\cdot m(\bm{x}),$
where $m(\bm{x})\vcentcolon=\min\left(\sigma_{\max}(\bm{x}),1-\|\sigma(\bm{x})\|_{2}^{2}\right)$.
@@
The upper bounds of Theorem 5.2 provide a tension. When the first term of $m(\bm{x})$ is maximized, the second is at its minimum, and vice-versa. In particular, the term that determines the upper bound depends on the uncertainty of the network’s output. When the network’s output is near random, e.g. at initialization, the first term will dominate with $\sigma_{\max}(\bm{x})\approx 1/C$. However, as the network becomes more certain with its prediction, the second term will start dominating: a more deterministic output $p(y\,|\,\bm{x})\rightarrow 1$ implies that $\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))\rightarrow 0$.
5.3 Empirical Verification: Classification
@@
Figure 1: MNIST FIM quantities for a 4-layer MLP utilizing sigmoid activation functions. Top: The estimated Fisher information, its variances, and the bounds of the variances across 4 parameter groups and 20 training epochs. The Fisher information (the thick green line) is estimated based on ^$\hat{\mathcal{I}}_{1}$ (^$\hat{\mathcal{I}}_{2}$ is almost identical and not shown for clarity). The standard deviation (square root of the variance) is shown for the variances and their bounds. Bottom: the log-ratio of the upper bound (UB) provided by Theorem 4.1 and the true variance. The closer to 0 the better quality the UB. On the right most column, the variance of ^$\hat{\mathcal{I}}_{2}$ vanishes: ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})=0\leq{\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$. Therefore related curves of ^$\hat{\mathcal{I}}_{2}$ are not shown.
@@
We experiment on the MNIST digit classification dataset (Deng, 2012) and a multilayer perceptron (MLP) consisting of 4 densely connected layers, sigmoid activation, and a dropout layer. The classification setting corresponds to using the categorical distribution with $C=10$ class labels. For a random $\bm{x}$ from the test set, we compute both estimators ^$\hat{\mathcal{I}}_{1}(\bm{\theta}\,|\,\bm{x})$ and ^$\hat{\mathcal{I}}_{2}(\bm{\theta}\,|\,\bm{x})$ using $N=5,000$ samples. We record the variances of each estimator and compute their bounds based on Theorem 4.1. The Fisher information of each individual parameter and its variances are aggregated based on the arithmetic average over each of 4 parameter groups (corresponding to the 4 layers) and examined throughout 20 training epochs.
@@
In Fig. 1, we present the variance scale of the estimators ^$\hat{\mathcal{I}}_{1}(\theta_{i}\,|\,\bm{x})$ and ^$\hat{\mathcal{I}}_{2}(\theta_{i}\,|\,\bm{x})$ in log-space; and the tightness of the bounds in Theorem 4.1 by consider the log-ratio $\log\frac{\textrm{UB}}{{\mathcal{V}}_{\mathrm{I}}(\theta_{i}\,|\,\bm{x})}$, where UB is the upper bounds in Theorem 4.1. In this experiment, the UB is much tighter than the lower bound (LB), which is omitted in the figures for better clarity. More experimental results are given in Appendix E.
@@
We varied the network architecture and the activation function. Across different settings, the proposed UB and LB are always valid. In Fig. 1, one can observe that the diagonal FIM and the associated variances have a small magnitude. For example, in the first layer, ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ and ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ are roughly $e^{-10}\approx 5\times 10^{-5}$. The log-ratio $\log\frac{\textrm{UB}}{{\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})}\approx 4$, meaning that UB is roughly 50 times larger than ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$. Comparatively, ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ has a tighter UB which is 20 times larger than itself. The UB serves as a useful hint on the order of magnitude of the variances. In Appendix C, we present tighter bounds which are more expensive to compute.
In the first 3 layers of the MLP, ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ presents a smaller value than ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$, meaning that ^$\hat{\mathcal{I}}_{1}$ can more accurately estimate the diagonal FIM. Interestingly, this is not true for the last layer: ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ becomes zero while ^$\hat{\mathcal{I}}_{1}$ presents the largest variance across all parameter groups and close to the scale of the Fisher information. Due to this, one should always prefer ^$\hat{\mathcal{I}}_{2}$ over ^$\hat{\mathcal{I}}_{1}$ for the last layer. In the last two layers, ^$\hat{\mathcal{I}}_{2}$ is in simple closed form without need of automatic differentiation (see Remark 4.4).
@@
The shape of the variance curves are sensitive to choices of the activation functions $\phi$ and inputs $\bm{x}$. In general, the first few epochs of the variance present more dynamics than the rest of the training process. If one uses log-sigmoid activations $\phi(t)=-\log(1+\exp(-t))$ (which is equivalent to $\phi(t)=-\textrm{SoftPlus}(-t)$, as per Remark 4.4), the variances of ^$\hat{\mathcal{I}}_{1}$ and ^$\hat{\mathcal{I}}_{2}$ only appear at the randomly initialized network and quickly vanish once training starts, as shown in Appendix E. In this case, the learner more easily approaches a nearly linear region of the loss landscape where the local optimum resides. In practice, one should estimate and examine the scale of variances — which should not be neglected as per Fig. 1 — before choosing a preferred diagonal FIM estimator for optimization and any other purposes.
@@ 6 Relationship with the “Empirical Fisher”
In some scenarios, calculating even the estimators ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ and ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$ of the diagonal FIM can still be expensive. Part of this cost comes from requiring label samples $\bm{y}_{k}$ for each $\bm{x}_{k}$ and their associated backward passes, Eqs. 3 and 4. For instance, when the FIM is used in an iterative optimization procedure, $\bm{y}_{k}$ need to be re-sampled at each learning step w.r.t. the current $\bm{h}$ and additional backpropagation needs to be run for each $k$.
As such, different ‘FIM-like’ objects have been explored which replace the samples from $p(\bm{y}\,|\,\bm{x})$ with samples taken from an underlying true (but unknown) data distribution $q(\bm{y}\,|\,\bm{x})$ Le Roux et al. (2007); Martens et al. (2010), i.e. utilizing samples from the data directly. We define the data’s joint distribution as $q(\bm{x},\bm{y})\vcentcolon=q(\bm{x})q(\bm{y}\,|\,\bm{x})$. In analogy to the FIM, the data Fisher information matrix (DFIM) can be defined as the PSD tensor $\mathrm{I}(\bm{\theta})\vcentcolon=\mathop{\mathbb{E}}_{q(\bm{x})}[\mathrm{I}(\bm{\theta}\,|\,\bm{x})]$, with
@@
where $\mathrm{I}(\bm{h}\,|\,\bm{x})=\mathop{\mathbb{E}}_{q(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})}\left[(\bm{t}(\bm{\hat{\mathrm{y}}})-\bm{\eta}(\bm{x}))(\bm{t}(\bm{\hat{\mathrm{y}}})-\bm{\eta}(\bm{x}))^{\top}\right]$ denotes the 2nd (non-central) moment of $(\bm{t}(\bm{\hat{\mathrm{y}}})-\bm{\eta}(\bm{x}))$ w.r.t. $q(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})$, and ${\partial\bm{h}}/{\partial\bm{\theta}}$ is the Jacobian of the map $\bm{\theta}\to\bm{h}$. In the special case that $q(\bm{y}\,|\,\bm{x})=p(\bm{y}\,|\,\bm{x},\,\bm{\theta})$, then $\mathrm{I}(\bm{\theta}\,|\,\bm{x})$ becomes exactly $\mathcal{I}(\bm{\theta}\,|\,\bm{x})$.
The DFIM $\mathrm{I}(\bm{\theta}\,|\,\bm{x})$ in Eq. 21 is a more general definition. Compared to the FIM $\mathcal{I}(\bm{\theta}\,|\,\bm{x})$, it yields a different PSD tensor on $\bm{\theta}$ parameter space (the neuromanifold) depending on a distribution $q(\bm{x},\,\bm{y})$, which is neither necessarily on the same neuromanifold nor necessarily parametric at all. Indeed, an asymmetry in the true data distribution and the empirical one induces a different geometric structure (Critchley et al., 1993). By definition, we have
$\mathrm{I}(\bm{\theta}\,|\,\bm{x})\succeq\left(\frac{\partial\mathrm{KL}}{\partial\bm{\theta}}\right)\left(\frac{\partial\mathrm{KL}}{\partial\bm{\theta}}\right)^{\top},$
where $\mathrm{KL}(\bm{\theta})\vcentcolon=\int q(\hat{\bm{y}}\,|\,\bm{x})\log\frac{q(\hat{\bm{y}}\,|\,\bm{x})}{p(\hat{\bm{y}}\,|\,\bm{x},\,\bm{\theta})}\,\mathrm{d}{\hat{\bm{y}}}$ is the Kullback-Leibler (KL) divergence, or the loss to be minimized. The DFIM can be regarded as a surrogate function of the squared gradient of the KL divergence. It is a symmetric covariant tensor and satisfies the same rule w.r.t. reparameterization as the FIM. Consider the reparameterization $\bm{\theta}\to\bm{\zeta}$, the DFIM becomes $\mathrm{I}(\bm{\zeta}\,|\,\bm{x})=\left({\partial\bm{\theta}}/{\partial\bm{\zeta}}\right)^{\top}\mathrm{I}(\bm{\theta}\,|\,\bm{x})\left({\partial\bm{\theta}}/{\partial\bm{\zeta}}\right)$.
Notice that $\hat{\bm{\eta}}(\bm{x})\vcentcolon=\mathop{\mathbb{E}}_{q(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})}[\bm{t}(\bm{\hat{\mathrm{y}}})]\neq\bm{\eta}(\bm{x})$ in general. As such, there will be a miss-match when utilizing $\mathrm{I}(\bm{h}\,|\,\bm{x})$ as a substitute for $\mathcal{I}(\bm{h}\,|\,\bm{x})$. However, as learning progresses and $p(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})$ becomes more similar to the data’s true labeling posterior $q(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})$, the DFIM will become closer to the FIM.
@@
where $(\bm{x}_{1},\bm{\hat{\mathrm{y}}}_{1}),\ldots,(\bm{x}_{N},\bm{\hat{\mathrm{y}}}_{N})$ are i.i.d. sampled from $q(\bm{x},\bm{\hat{\mathrm{y}}})$. Similarly to ^$\hat{\mathcal{I}}_{1}(\theta_{i})$, an estimator with a fixed input $\bm{x}$ can be considered, denoted as $\hat{\mathrm{I}}(\theta_{i}\,|\,\bm{x})$.
Given the computational benefits of being able to use data directly (and not require a separate sampling routine), many popular optimization methods utilize (an approximation of) the empirical Fisher. For instance, in the Adam optimizer Kingma & Ba (2015), the empirical Fisher is used to obtain an approximation of the diagonal FIM and to adapt to the geometry of the data. However, the change of switching from samples $\bm{y}_{k}$ to $\bm{\hat{\mathrm{y}}}_{k}$ is anything but superficial (Martens, 2020, Chapter 11)$\hat{\mathrm{I}}(\bm{\theta})$ is not an unbiased estimator of $\mathcal{I}(\bm{\theta})$ as $\mathrm{I}(\bm{h}\,|\,\bm{x})$ is different from $\mathcal{I}(\bm{h}\,|\,\bm{x})$.
The biased nature of the empirical Fisher affects the other moments of the estimator as well. In particular, we do not have the same equivalence of covariance and the metric being pulled back by $\bm{\theta}\rightarrow\bm{h}$ Sun (2020).
@@
As a result, although the estimator $\hat{\mathrm{I}}(\theta_{i}\,|\,\bm{x})$ emits a similar form to ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ (i.e., Eq. 11), the bounds are expressed via non-central moments of $(\bm{t}(\bm{\hat{\mathrm{y}}})-\bm{\eta}(\bm{x}))$ and not the central moments of $\bm{t}$ which were used in the estimators of the FIM — noting the miss-match in $\hat{\bm{\eta}}(\bm{x})\neq\bm{\eta}(\bm{x})$. Some corresponding DFIM bounds are characterized in Appendix F.

### v3

@@ 5 Case Studies
Suppose that Eq. 1 is a categorical distribution. With $\sigma_{\max}(\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\max_{k}\sigma_{k}(\bm{x})$:
$\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))\leq m(\bm{x});\quad\textrm{and}\quad\tilde{\lambda}_{\max}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}(\bm{h}\,|\,\bm{x})\otimes\mathcal{I}(\bm{h}\,|\,\bm{x}))\leq 2\cdot m(\bm{x}),$
where $m(\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\min\left(\sigma_{\max}(\bm{x}),1-\|\sigma(\bm{x})\|_{2}^{2}\right)$.
@@
This upper bounds provides a tension. When the first term of $m(\bm{x})$ is maximized, the second is minimized, and vice-versa. In particular, the dominating term depends on the uncertainty of the NN’s output. When the NN’s output is near random, e.g. at initialization, the first term will dominate with $\sigma_{\max}(\bm{x})\approx 1/C$. However, as the NN becomes more certain with its prediction, the second term will start dominating: a more deterministic output $p(y\,|\,\bm{x})\rightarrow 1$ implies that $\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))\rightarrow 0$.
@@
Figure 2: MNIST for a 4-layer MLP with sigmoid activations. Top: The estimated Fisher information (FI), variances, and variance bounds across 4 parameter groups and 20 training epochs. The FI (green line) is estimated using ^$\hat{\mathcal{I}}_{1}$ (^$\hat{\mathcal{I}}_{2}$ is almost identical and not shown for clarity). The s.t.d. (square root of variance) is shown for variances and their bounds. Bottom: the log-ratio of Theorem 4.1’s upper bounds (UBs) and the true variances. The closer to 0, the better the UB. In the right most column, the variance of ^$\hat{\mathcal{I}}_{2}$ vanishes: ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})=0\leq{\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$. Thus related curves of ^$\hat{\mathcal{I}}_{2}$ are not shown.
Empirical Verification: Classification
@@
We examine the MNIST classification task (LeCun et al., 2010) (CC BY-SA 3.0) using multilayer perceptrons (MLP) with four densely connected layers, sigmoid activations, and a dropout layer. For classification, we consider a categorical distribution with $C=10$ class labels. For a random $\bm{x}$ from the test set, we compute both estimators ^$\hat{\mathcal{I}}_{1}(\theta_{i}\,|\,\bm{x})$ and ^$\hat{\mathcal{I}}_{2}(\theta_{i}\,|\,\bm{x})$ using $N=5,000$ samples. We record the variances of each estimator and compute their bounds based on Theorem 4.1. For all 20 training epochs, the Fisher information (FI) and their variances of individual parameters are aggregated via arithmetic averages over four parameter groups (corresponding to the four layers).
@@
In Fig. 2, we present the variance scale of the estimators ^$\hat{\mathcal{I}}_{1}(\theta_{i}\,|\,\bm{x})$ and ^$\hat{\mathcal{I}}_{2}(\theta_{i}\,|\,\bm{x})$ in log-space; and the tightness of the bounds in Theorem 4.1 by consider the log-ratio $\log\frac{\textrm{UB}}{{\mathcal{V}}_{\mathrm{I}}(\theta_{i}\,|\,\bm{x})}$, where UB is the upper bounds in Theorem 4.1. In this experiment, the UB is much tighter than the lower bound (LB), which is omitted in the figures for clarity. More experimental results are given in Appendix F.
@@
We varied the NN’s architecture and activation function. Across different settings, the proposed UB and LB are always valid. In Fig. 2, one can observe that the diagonal FIM and the associated variances have a small magnitude. For example, in the first layer, ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ and ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ are roughly $e^{-10}\approx 5\times 10^{-5}$. The log-ratio $\log\frac{\textrm{UB}}{{\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})}\approx 4$ means that the UB is roughly 50 times larger than ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$. Comparatively, ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ has a tighter UB which is approximately 10 times larger than itself. The UB serves as a useful hint on the order of magnitude of the variances. In Appendix D, we present tighter bounds which are more expensive to compute.
@@
In the first three layers of the MLP, ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ presents a smaller value than ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$, meaning that ^$\hat{\mathcal{I}}_{1}$ can more accurately estimate the diagonal FIM. Interestingly, this is not true for the last layer: ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ becomes zero while ^$\hat{\mathcal{I}}_{1}$ presents the largest variance across all parameter groups. Due to this, one should always prefer ^$\hat{\mathcal{I}}_{2}$ over ^$\hat{\mathcal{I}}_{1}$ for the last layer. In the last two layers, ^$\hat{\mathcal{I}}_{2}$ is in simple closed form and, hence, does not need automatic differentiation to calculate (see Remarks 4.4 and 4.5). The shape of the variance curves are sensitive to the choices of activation functions $\phi$ and inputs $\bm{x}$. In general, the variance in the first few epochs presents more dynamics than the rest of the training process. If one uses log-sigmoid activations $\phi(t)=-\log(1+\exp(-t))$ (which is equivalent to $\phi(t)=-\textrm{SoftPlus}(-t)$, as per Remark 4.4), the variances of ^$\hat{\mathcal{I}}_{1}$ and ^$\hat{\mathcal{I}}_{2}$ only appear in the randomly initialized NN and quickly vanish once training starts, as shown in Appendix F. In this case, the learner more easily approaches a nearly linear region of the loss landscape where local optima lie. In practice, one should estimate and examine the scale of variances — which should not be neglected as per Fig. 2 — before choosing a preferred diagonal FIM estimator.
@@ 6 Relationship with the “Empirical Fisher”
In some scenarios, even the estimators of the diagonal FIM ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ and ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$ can be prohibitively expensive. Part of the cost comes from requiring label samples $\bm{y}_{k}$ for each $\bm{x}_{k}$, as per Eq. 3. For example, when the FIM is used in an iterative optimization procedure, $\bm{y}_{k}$’s need to be re-sampled at each learning step w.r.t. the current $\bm{h}$ alongside their backpropagation (accounting for sampling).
As such, alternative ‘FIM-like’ objects have been explored which replace the samples from $p(\bm{y}\,|\,\bm{x})$ with samples from an underlying true (but unknown) data distribution $q(\bm{y}\,|\,\bm{x})$ Le Roux et al. (2007); Martens et al. (2010). We define the data’s joint distribution as $q(\bm{x},\bm{y})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}q(\bm{x})q(\bm{y}\,|\,\bm{x})$. Analogous to the FIM, the data Fisher information matrix (DFIM) can be defined as the PSD tensor $\mathrm{I}(\bm{\theta})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathop{\mathbb{E}}_{q(\bm{x})}[\mathrm{I}(\bm{\theta}\,|\,\bm{x})]$, with
@@
where $\mathrm{I}(\bm{h}\,|\,\bm{x})=\mathop{\mathbb{E}}_{q(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})}\left[(\bm{t}(\bm{\hat{\mathrm{y}}})-\bm{\eta}(\bm{x}))(\bm{t}(\bm{\hat{\mathrm{y}}})-\bm{\eta}(\bm{x}))^{\top}\right]$ denotes the 2nd (non-central) moment of $(\bm{t}(\bm{\hat{\mathrm{y}}})-\bm{\eta}(\bm{x}))$ w.r.t. $q(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})$, and ${\partial\bm{h}}/{\partial\bm{\theta}}$ is the Jacobian of the map $\bm{\theta}\to\bm{h}$. In the special case that $q(\bm{y}\,|\,\bm{x})=p(\bm{y}\,|\,\bm{x};\bm{\theta})$, then $\mathrm{I}(\bm{\theta}\,|\,\bm{x})$ becomes exactly $\mathcal{I}(\bm{\theta}\,|\,\bm{x})$.
The DFIM $\mathrm{I}(\bm{\theta}\,|\,\bm{x})$ in Eq. 17 is a more general definition. Compared to the FIM $\mathcal{I}(\bm{\theta}\,|\,\bm{x})$, it yields a different PSD tensor on the $\bm{\theta}$ parameter space (the neuromanifold) depending on a distribution $q(\bm{x},\bm{y})$, which is neither necessarily on the same neuromanifold nor necessarily parametric at all. The asymmetry in the true data distribution and the empirical one results in different geometric structures (Critchley et al., 1993). By definition, we have $\mathrm{I}(\bm{\theta}\,|\,\bm{x})\succeq\left({\partial\mathrm{KL}}/{\partial\bm{\theta}}\right)\left({\partial\mathrm{KL}}/{\partial\bm{\theta}}\right)^{\top}$, where $\mathrm{KL}(\bm{\theta})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\int q(\hat{\bm{y}}\,|\,\bm{x})\log\tfrac{q(\hat{\bm{y}}\,|\,\bm{x})}{p(\hat{\bm{y}}\,|\,\bm{x};\bm{\theta})}\,\mathrm{d}{\hat{\bm{y}}}$ is the Kullback-Leibler (KL) divergence, or the loss in a parameter learning scenario. The DFIM can be regarded as a surrogate function of the squared gradient of the KL divergence. It is a symmetric covariant tensor and satisfies the same rule w.r.t. reparameterization as the FIM. Consider the reparameterization $\bm{\theta}\to\bm{\zeta}$, the DFIM becomes $\mathrm{I}(\bm{\zeta}\,|\,\bm{x})=\left({\partial\bm{\theta}}/{\partial\bm{\zeta}}\right)^{\top}\mathrm{I}(\bm{\theta}\,|\,\bm{x})\left({\partial\bm{\theta}}/{\partial\bm{\zeta}}\right)$.
Notice that $\hat{\bm{\eta}}(\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathop{\mathbb{E}}_{q(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})}[\bm{t}(\bm{\hat{\mathrm{y}}})]\neq\bm{\eta}(\bm{x})$ in general. As such, there will be a miss-match when utilizing $\mathrm{I}(\bm{h}\,|\,\bm{x})$ as a substitute for $\mathcal{I}(\bm{h}\,|\,\bm{x})$. However, as learning progresses and $p(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})$ becomes more similar to the data’s true labeling posterior $q(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})$, the DFIM will become closer to the FIM.
@@
where $(\bm{x}_{1},\bm{\hat{\mathrm{y}}}_{1}),\ldots,(\bm{x}_{N},\bm{\hat{\mathrm{y}}}_{N})$ are i.i.d. sampled from $q(\bm{x},\bm{\hat{\mathrm{y}}})$. Similar to ^$\hat{\mathcal{I}}_{1}(\theta_{i}\,|\,\bm{x})$, an estimator with a fixed input $\bm{x}$ can be considered, denoted as $\hat{\mathrm{I}}(\theta_{i}\,|\,\bm{x})$.
Given the computational benefits of using the data directly — bypassing a separate sampling routine — many popular optimization methods employ the empirical Fisher or its approximation. For instance, the Adam optimizer Kingma and Ba (2015) uses the empirical Fisher to approximate the diagonal FIM. However, switching from sampling $\bm{y}_{k}$ to $\bm{\hat{\mathrm{y}}}_{k}$ is anything but superficial (Martens, 2020, Chapter 11)$\hat{\mathrm{I}}(\bm{\theta})$ is not an unbiased estimator of $\mathcal{I}(\bm{\theta})$ as $\mathrm{I}(\bm{h}\,|\,\bm{x})$ is different from $\mathcal{I}(\bm{h}\,|\,\bm{x})$.
The biased nature of the empirical Fisher affects the other moments as well. In particular, we do not have the same equivalence of covariance and the metric being pulled back by $\bm{\theta}\rightarrow\bm{h}$ Sun (2020).
@@
As a result, although the variance of the estimator $\hat{\mathrm{I}}(\theta_{i}\,|\,\bm{x})$ takes a similar form to ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ (i.e., Eq. 8), its sufficient statistic terms do not exclusively consist of central moments. Noting the miss-match in $\hat{\bm{\eta}}(\bm{x})\neq\bm{\eta}(\bm{x})$, Lemma 6.1 reveals an additional term which shifts $\mathrm{I}(\bm{h}\,|\,\bm{x})$ away from the 2nd central moment of $\bm{t}(\bm{\hat{\mathrm{y}}})$ (w.r.t. $q(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})$). Instead, these sufficient statistic terms correspond to non-central moments of $\bm{t}(\bm{\hat{\mathrm{y}}})-\bm{\eta}(\bm{x})$. Some corresponding empirical Fisher / DFIM bounds are characterized in Appendix G.